FEL Upgrade Project Weekly Brief
November 6 – 10, 2000


We had a very successful week of FEL operations producing a new average power record for sub-picosecond class 1 micron light by generating over 340 watts with third harmonic lasing.  We also attempted our first experiments with doubling (to green) and quadrupling (to UV) the 1 micron light achieving 28% and 5% conversion efficiencies on our initial trails with non-optimized non-linear conversion optics.


We prepared for submission next week three proposals in response to the Joint Technology Office call for proposals on high energy laser technology.  Our response topics were:  (1) femto second laser damage effects; (2) high average current beam transport; and (3) high average current injector development.

On Thursday we had a visit from Elizabeth Turpen, legislative staff member from Senator Domenici’s office who deals with defense issues.  We also had a visit from NRL, DARPA and Lockheed Martin personnel who are interested in discussing possible lithography activities with the synchrotron which is being considered for addition to the FEL Facility.

WBS 3 (Beam Physics)

Evaluation of Upgrade injector dynamics continues, with a 135 pC solution using rectangular dipoles established with marginally adequate performance.

WBS 4 (Injector)

The second nitrogen implanted electrode for the field emission test system was run at 24 MV/m (5 mm gap at 120 kV) and had an average field emission current of 100pA.  This initial data is better than the first implanted electrode which showed 145 pA at 24 MV/m.

WBS 6 (RF)

In zone 4, the problems with interlocks were nearly all solved.  A few cross-connects still need to be resolved and 2 test heads in the vault need to be replaced when the run is over.

No work was accomplished in zone 3.

WBS 9 (Transport)

o We sent out preliminary dipole drawings with a request for a budgetary price to six vendors.
o We continued detailing the GW dipole and started procurement of its steel.
o The GW 3D model is working, studies started to determine the effectiveness of a Purcell Gap and the choice of materials.
o The effective length and K1 were determined from the existing baseline GX model.  K1 was only half of what was found by 2D models run by J. Karn in April.  An error was found in how K1 was calculated in those 2D models, and once corrected, agreed with the 3D models.  However, the errant values provided by J. Karn were used to optimize the existing lattice.  D. Douglas has taken a preliminary look at the difference and does not anticipate any critical implications.
o 3 inch Quad
o Prototype Fabrication: Inspection of the latest aluminum sample quadrant showed conflicting results between CMM and optical tooling measurements.  The two techniques showed disagreements between 0.005" and 0.010".
o Rather than spending an endless amount of time perfecting a single quadrant the decision was made to concentrate on building an assembled core.  The shop was instructed to machine the four steel quadrants at best effort. An inspection will be performed of the assembled quadrants and any required "corrective machining" will be performed on the assembly.  This will get us a prototype sooner to qualify the design
o In parallel we will pursue manufacturing options for the production magnets.  Various components of the quad package have been sent to five additional vendors for budgetary quotes.
o Measurement Probe: Glass material for the probe has arrived, drawings of the probe have started, and printed circuit boards for the coils are being fabricated.
o Debbie Hedrick of the vacuum group joined the coordination effort, firming up the requirements for vacuum part orders.

WBS 10 (Wiggler)

o Drawings for the manifold mounts were revised to accommodate mounting the wiggler cover and are in procurement.
o Drawings for the buss bar mounting were completed and are also in procurement.
o No assembly work was done because the technician is on vacation

WBS 11 (Optics)

The drawings for the mirror test stand underwent final check and are now being printed for sign-off.

We have ordered a gimbal mount that will be used in the test stand, and is being considered for the optical cavity as well.  We will use it in our design of the optical cavity, so we can design a vacuum vessel for it, and feed that information back to WBS 3 in order for them to more quickly finalize the lattice.

We have ordered additional pyroelectric detectors to provide additional diagnostic capabilities.


This was a big week for us.  Using a new outcoupler optimized for 1 micron operation, we measured 220 W of power at 1028 nm upstairs, and 340 W downstairs, almost immediately after the outcoupler.  These results solidly place us in record-making levels of power for an ultrafast laser.  The ratio 65%, agrees with a calculation based on the literature values of silver's reflectivity.

We took this light (while in pulsed mode), and doubled it in lithium iodate, measuring a conversion efficiency of ~ 28% to ~ 530 nm.  This efficiency dropped considerably when we attempted to double while running cw, due to crystal heating.  Then, working with W. Cooke (College of William and Mary), we tripled and quadrupled the ~ 1 micron light.  The quadrupling conversion efficiency was actually fairly reasonable, ~ 20% from the SHG light, or ~ 5% from the ~1 micron light.  The third harmonic generation conversion efficiency was poor, due to temporal dispersion effects in the material.  We're obtaining more optimized materials, but do not expect them until later this month.  All this work is important, both for our current users wishing to do micromachining, or AMO physics.

We also delivered beam to a group from Kansai Research Institute and RPI, here to generate high intensities of THz radiation.  To date, they have been very successful, measuring both relatively high intensities, and the duration of the radiation.

Privacy and Security Notice
Privacy and Security Notice